OpenStudy (anonymous):

-7x^2=-448 taking the square root and completing the square please show steps

OpenStudy (anonymous):

you do not have to complete the square for this one divide both sides by \(-7\) then take the square root of both sides don't forget the \(\pm\)

OpenStudy (max1409):

https://www.mathway.com

pooja195 (pooja195):

\[\huge~\rm~-448\div-7=?\]

OpenStudy (anonymous):

64

geerky42 (geerky42):

Yep, so now you have \(x^2 = 64\) Your next step is to take square root of both sides.

pooja195 (pooja195):

\[\huge~\rm~\sqrt{x^2}=\sqrt{64}\]

OpenStudy (anonymous):

64 is 8

pooja195 (pooja195):

good and the square root of X^2 ?

OpenStudy (anonymous):

x

pooja195 (pooja195):

ok so whats the answer?

OpenStudy (anonymous):

8x

OpenStudy (anonymous):

so the equation for it would be like the steps

pooja195 (pooja195):

Nope remeber the = sign it remains there! \[\huge~\rm~x=\pm8\]

OpenStudy (anonymous):

what would be the steps

pooja195 (pooja195):

We just did it ... we need to divide both sides by -7 that way we can isolate the x \[\huge~\rm~\frac{ -7x^2=-448 }{ 7 } \] We were left with \[\huge~\rm~x^2=64\] we need to solve for x so we need to square root both of those \[\huge~\rm~\sqrt{x^2}=\sqrt{64}\]

pooja195 (pooja195):

And when we did this we ended up with... \[\huge~\rm~x=\pm8\]

pooja195 (pooja195):

understood ?

OpenStudy (anonymous):

yes i think so if i turned these steps it would be the correct formulas and stuff?

pooja195 (pooja195):

Can you please rephrase that ?

OpenStudy (anonymous):

i'm suppose to show my work when i turn it in so if i turn these steps in would it be the correct formula?

OpenStudy (anonymous):

can you help me with one more?

pooja195 (pooja195):

._. this is correct but i expect u to be able to do the next one by yourself if it is like this...i can look over the work :)

OpenStudy (anonymous):

ok well the next one is a little different x^2-10x+22=-2

OpenStudy (anonymous):

do i do the same thing?

pooja195 (pooja195):

We need to get it into standered for first \[\huge~\rm~ax^2+bx+c=0\]

OpenStudy (anonymous):

fill it in with the number from the equation?

pooja195 (pooja195):

x^2-10x+22=-2 we add the -2 to the - \[\huge~\rm~22-2=?\]

OpenStudy (anonymous):

well -24

pooja195 (pooja195):

my mistake sorry it would be just 24

pooja195 (pooja195):

\[\huge~\rm~x^2-10x+24=0\]

pooja195 (pooja195):

what numbers multiply to make 24 but adds to make -10

OpenStudy (anonymous):

4 an 6 im not sure

pooja195 (pooja195):

good -4 and -6 then (x-6) (x-4) set up 2 diffrent equations x-6=0 <---solve for x x-4=0 <----solve for x

OpenStudy (anonymous):

x=6 and x=4

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

yep

OpenStudy (anonymous):

and then?

pooja195 (pooja195):

thats it

OpenStudy (anonymous):

o so the steps are ?

OpenStudy (anonymous):

i got to go